1. At what month (m) will the bill be the same? Show all of your work using the line paper or in your physical notebook . Hint: Set both expressions equal to each other and solve for m The latest Android will cost you a down payment of $110 and \$70/mont with T-Mobile. The latest iPhone will cost you a down payment of $120 and \$65/month with T-Mobile.


1
Expert's answer
2021-06-27T16:10:55-0400

Assume that m is the month when the bills will be equal. So, equating two bills with each other we can make the equation:

   
"110+70 m = 120+65 m""70 m - 65 m = 120 - 110""5 m = 10 , m=10\/5=2"